OpenStudy (ksmd13):

If x^2 = 30, what is the value of x? ±60 ±15 ±square root of 30 ±square root of 15

OpenStudy (xguardians):

What do you think the first step is?

OpenStudy (anonymous):

what is your opinion please ?

OpenStudy (ksmd13):

well 15 x 2 = 30 so it would be b or d right?

OpenStudy (ksmd13):

idk what to do next

OpenStudy (ksmd13):

@xGuardians @jhonyy9

OpenStudy (anonymous):

x^2 = 30 make squerrot on both sides and so what will get ?

OpenStudy (anonymous):

squarroot -- sorry

OpenStudy (anonymous):

(x^2)^(1/2) = 30^(1/2) yes ?

OpenStudy (anonymous):

but you know that sqrtx^2 = +/- x yes ?

OpenStudy (anonymous):

so in the othjer side too in case of 30

OpenStudy (anonymous):

other side - sorry

OpenStudy (ksmd13):

i'm really confused sorry

OpenStudy (ksmd13):

so i multiply ^1/2 on both sides?

OpenStudy (anonymous):

(x^2)^(1/2) = x^(2*(1/2)) = x^(2/2 = x yes ? no this expponent of 1/2 meaning the squareroot

OpenStudy (anonymous):

bc sqrt(x) = x^(1/2) you dont know it ?

OpenStudy (ksmd13):

oh ok

OpenStudy (anonymous):

so this mean that you understand it now sure ?

OpenStudy (anonymous):

ok was my pleasure good luck bye bye